OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 9 white balls and 9 yellow balls and the second urn contains 8 white balls and 3 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

They're independent events, so you just multiply the probabilities for the individual urns.

OpenStudy (anonymous):

urns are supposed to contain ashes, not balls wtf!

OpenStudy (anonymous):

lol huuh ??

OpenStudy (anonymous):

what do i multiple

OpenStudy (anonymous):

the probability of a white ball from the first urn and the probability of a while ball from the second urn

OpenStudy (anonymous):

thanks

Directrix (directrix):

The events are independent. The outcome of the first draw does not affect the outcome of the second draw. ---- P(white1 and white2) = P(white1)*P(white2) = (9/18)*(8/11) = 4/11. @KeemoD143